# DualFluidNet: an Attention-based Dual-pipeline Network for Accurate and Generalizable Fluid-solid Coupled Simulation

### v1

@@ 5 Conclusion
We have designed an Attention-based Dual-pipeline network architecture, which provides stable global fluid control and ensures adherence to physical law constraints in fluid simulation. The Dual-pipeline feature fusion architecture also holds the potential to inspire work in other domains. Additionally, we propose a Type-aware Input Module, which plays a pivotal role in facilitating the fluid-solid particle coupling. Our network represents a step forward in utilizing neural networks to simulate physically compliant fluids, enhancing the potential for real-time simulation generalization across diverse fluid scenarios, particularly in scenarios such as interactive gaming environments, and virtual reality simulations. There remains untapped potential in enabling neural networks to learn real-world physics. We will make our code publicly available to contribute to the ongoing research.

### v2

@@ 5 Conclusion
We have designed an Attention-based Dual-pipeline network architecture, which provides stable global fluid control and ensures adherence to physical law constraints in fluid simulation. The Dual-pipeline feature fusion architecture also holds the potential to inspire work in other domains. Additionally, we propose a Type-aware Input Module, which plays a pivotal role in facilitating the fluid-solid particle coupling. Our network achieves not merely a quantitative improvement in performance metrics, but a qualitative leap in neural network-based simulation fidelity adhering to the laws of the physical world, enhancing the potential for real-time simulation generalization across diverse fluid scenarios, particularly in scenarios such as interactive gaming environments, and virtual reality simulations. However, there still remains untapped potential in enabling neural networks to learn real-world physics. We will make our code publicly available to contribute to the ongoing research.
Acknowledgments
The work reported in this paper is supported by the National Natural Science Foundation of China [52375514].
CRediT authorship contribution statement
Yu Chen: Methodology, Algorithm, Validation, Writing - original draft. Shuai Zheng: Conceptualization, Funding acquisition, Project administration, Supervision, Writing - review & editing. Menglong Jin: Visualization, Data curation. Yan Chang: Visualization. Nianyi Wang: Investigation, Writing - review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
We have shared the link to our code and data on the following GitHub repository: https://github.com/chenyu-xjtu/DualFluidNet.